# Automated Quantum Algorithm Synthesis

### v2

@@ I Introduction
Quantum computers promise exponential speed-ups, yet there are only a handful of classes of quantum algorithms that realise this. Shor Shor (2003) provides us with two possible explanations. Firstly, quantum computers operate so differently from classical computers that our existing techniques for algorithm design and our intuitions about computation may no longer be effective. Secondly, it is possible that only a limited number of such quantum algorithms exist and that many or all of them may have already been discovered.
@@
Given the steadily growing Quantum Algorithm Zoo  [^1] https://quantumalgorithmzoo.org/ we might retain some optimism with regards to the latter point. Nevertheless, developing novel protocols—especially those that generate new classes of quantum algorithms, of which Shor’s factorising algorithm Shor (1999) is an example—is clearly a difficult task. Computational methods could offer a way to bridge the intuition gap and facilitate the discovery process. We propose such a method that uses a bottom-up approach, where a collection of well-constructed examples is translated into an algorithm.
@@
Our approach follows the methodology of neural architecture search (NAS) where the task of automated design is divided into three parts, search space, search strategy and performance estimation strategy Elsken et al. (2019); Wistuba et al. (2019); Liu et al. (2022); White et al. (2023). Neural architecture search strategies are widely used in machine learning to improve performance in computer vision, natural language processing and time series forecasting Liu et al. (2022). More recently, NAS methods have also been applied in the context of variational quantum algorithms, for a recent survey see Martyniuk et al. (2024).
@@
Quantum algorithms, which are associated with highly structured circuits, tend to be modular and to contain repeating patterns. To exploit these facts we use a domain-specific language (DSL)—specific syntax and rules for compactly expressing quantum circuits via modular building blocks, implemented as an open source Python package Lourens et al. (2023); Lourens (2024a, b). In this DSL, quantum circuits are abstracted beyond the usual gate-sequence description and scale automatically with problem size. This allows us to optimise the structure of the algorithm, as opposed to the quantum circuit at the level of individual gates. Previous work on automated quantum algorithm design Almuqbil and Felemban (2024); Lin et al. (2020); Wan et al. (2018); Potoček et al. (2018); Bang et al. (2014); Bang and Yoo (2014); Hutsell and Greenwood (2007); Surkan and Khuskivadze (2002) was limited to gate-based circuit implementations of algorithms for small circuit sizes, usually $2$-$8$ qubits. In contrast, our learnt algorithms automatically generalise to arbitrary problem size which has not been demonstrated previously. Remarkably, this was achieved by using training examples related to just three small circuit sizes, namely $2$, $3$ and $4$ qubits or $3$, $4$ and $5$ qubits.
@@
Our method is compatible with any gradient-free optimisation procedure, specifically, we use evolutionary search DE (1989); Schwefel and Rudolph (1995); Rechenberg (1973). The performance of candidate algorithms is measured by how closely their behaviour matches the examples provided in the training set. In addition, the evaluation takes into account the uniformity of the performance across all examples and the complexity of the candidate algorithm. We demonstrate our methodology by rediscovering three well-known quantum algorithms, the quantum Fourier transform, the Deutsch-Jozsa algorithm and Grover’s search. This shows that our method works for oracle-based problems and variational circuits. Our approach also has the necessary functionality for designing both measurement-based algorithms and ansatzes for variational quantum circuits Lourens et al. (2023, 2025).
{minted} python deutsch_jozsa = ( QCycle(H) + QPivot(Z, ”*1”) + Oracle + QPivot(Z, ”*1”) + QCycle(H) ) * r
{minted} python grover = ( Oracle + QCycle(H) + QCycle(X) + QPivot(H, ”*1”) + QPivot(MCX, ”*1”) + QPivot(H, ”*1”) + QCycle(X) + QCycle(H) ) * r
{minted} python qft = ( QPivot(H, ”1*”) + QPivot(CRphi, ”1*”) + Qmask(”1*”) ) * r
@@ II Results
Using our domain-specific language combined with evolutionary search we were able to rediscover three well-known quantum algorithms, the quantum Fourier transform (QFT), the Deutsch-Jozsa algorithm and Grover’s search. Figure 1 shows two equivalent $n$-qubit representations of three automatically designed algorithms, corresponding to each of the learning tasks. On the left-hand side of Fig. 1 are the leant algorithms expressed in the DSL. On the right-hand side are the corresponding quantum circuit representations, where each circuit matches the common, hardware-agnostic, implementation of each algorithm Nielsen and Chuang (2001). Fig. 1 illustrates how the DSL compactly describes the general $n$-qubit implementation of the algorithms. This allows us to immediately interpret the algorithm structure, as well as identify symmetries and patterns. For example, the symmetric structure of the diffusion term in Grover’s search is immediately identifiable.
An evolutionary algorithm is used to explore the search space where new candidate algorithms are generated by changing or combining the fittest candidates from a population of algorithms. Remarkably, we found that to converge to the general algorithm, it was sufficient to evaluate each candidate on three small circuit sizes, namely $2$, $3$ and $4$ qubits for the Deutsch-Jozsa algorithm and QFT, and $3$, $4$ and $5$ qubits for Grover’s search. A candidate algorithm has a high fitness if it executes the target algorithm correctly and has low complexity, discussed further in section IV.3. Furthermore, we did not tune the hyper-parameters of the evolutionary search. These were kept as consistent as possible across all experiments. That the same hyper-parameters were good enough for all three tasks, indicates that our search strategy is stable and reliable.
@@
The right panel of Fig. 2 shows the changing composition of the population. The slope of the contour lines indicates the rate of displacement of the poorest performing candidates from the top $200$ algorithms. The steeper the slope of a contour line, the slower poorer-performing candidates are displaced. The general trend is that the gradient of each contour line decreases over time as the current best algorithm, its duplicates and related variants, propagate rapidly and take up an increasing proportion of the population.
Fig. 3 shows the evolutionary path, represented as a directed graph, that led to Grover’s search. The graph contains a roughly equal distribution of mutation types, represented by different edges. This indicates that all four mutation types played an important role in producing the final algorithm, which was the case for all experiments across all three tasks. In the literature, it has been suggested that mutations that combine two candidates, in our case the ‘chop-and-join’ mutation, do not lead to better-performing architectures White et al. (2023); Rattew et al. (2020). Our results indicate that in the context of algorithm design, this is not the case. A possible explanation is that we break each architecture into two pieces before recombining them, rather than simply appending algorithms head-to-tail. This provides an opportunity for useful sections of each algorithm to be combined. Dividing each algorithm into two randomly determined pieces before recombination also ensures that the size of candidates does not double with each application of this mutation. Furthermore, random division ensures diversity in the size of candidates.
@@
We demonstrate the robustness of our method by investigating the effect of increasing search space size. The DSL that was used to represent our search space is implemented as an open-source Python package called Hierarqcal Lourens et al. (2023, 2025). See Section IV.1 and Fig. 1 for an illustration of the basic functionality of Hierarqcal. In this framework, algorithms are built up of motifs that encode different gate placement patterns. Each motif has several properties, each of which can take on a possibly infinite number of discrete values. Motif properties include mapping, step, stride, offset, boundary condition, global and merge within patterns, weight sharing and edge order. Selecting appropriate sets of discrete values for each of these properties determines the configuration space, which in turn determines the set of possible algorithms that can be constructed, i.e. the search space.
@@
In Fig. 4 we see that the number of candidates evaluated in each experiment was a small fraction of the size of the search space, indicating the efficiency of our evolutionary strategy. The number of evaluations required grows with the size of the search space, though favourable at a decreasing rate. A total of $50$ experiments were performed for each combination of task and search space size, and successful runs are shown in Fig. 4. An experiment is considered successful if it is completed within a certain runtime limit and if certain stopping conditions are met, which we describe in detail in Section IV.3. Table 1 shows the approximate success rates for each task and notes the runtime limits in core processing unit (CPU) hours. Taking the success rates into account, our method is robust and can plausibly be run on a personal computer – $6$-$14$ CPUs – with a $99\%$ chance of success within a single day in most cases.
@@
[…]
@@ III Discussion
Our study demonstrates that computational methods can effectively be used to discover quantum algorithms. We successfully rediscover the quantum Fourier transform (QFT), the Deutsch-Jozsa algorithm, and Grover’s search validating our approach. Central to our method is a domain-specific language, which captures the inherent scalability of algorithms. Further, the DSL biases the algorithms in the search space to be structured, capturing modularity and repeating patterns. As a result, the learnt algorithms are immediately interpretable and analysable. Notably, convergence to the general algorithm requires evaluation on only three small circuit sizes, containing at most $5$ qubits, which significantly reduces the computational cost.
@@
Furthermore, the DSL contains a small number of building blocks, that in turn depend on a limited number of properties. Therefore, we can greatly reduce the size of the search space compared to using a gate-based circuit description. Importantly, we are able to reduce the search space size without sacrificing expressivity, since the DSL captures the essential characteristics of algorithms, namely structure and scalability.
@@
By using a bottom-up approach, we reformulate the task of algorithm design as a task of creating a set of well-constructed training examples. Our performance estimation strategy then evaluates candidate algorithms based on their ability to generalise across these training examples. We prioritize simpler algorithms by incorporating a complexity penalty while simultaneously rewarding algorithms that perform uniformly across training examples. This, combined with the attributes mentioned above, enables us to automatically design the $n$-qubit realisation of a quantum algorithm.
@@ IV Methods
Much of our success is due to the use of a domain-specific language. It allows us to learn the general algorithm structure rather than a specific unitary implementation of the algorithm for some number of qubits. The DSL was adapted from its initial formalism detailed in Lourens et al. (2023), where many of the core ideas were introduced. For a recent overview, see the blog post Lourens (2024a), and for the associated Python package Hierarqcal, see the GitHub repository Lourens (2024b).
The original inspiration for the DSL came from the field of neural architecture search Liu et al. (2018). In line with NAS we divide our automated algorithm design approach into three components: search space, search strategy, and performance estimationElsken et al. (2019); Wistuba et al. (2019); Liu et al. (2022). The DSL is used to specify the search space, see Section IV.1. Owing to the modular nature of our DSL, evolutionary search is a natural choice for search strategy, detailed in Section IV.2. The performance estimation method used to evaluate candidate algorithms is described in Section IV.3. Finally, in Section IV.4, we provide an overview of each design task, with a particular focus on problem setup.
@@ IV.1 Search space
The motifs are the building blocks of our DSL and are used to generate hypergraphs associated with a unitary operation. The nodes of the hypergraph correspond to qubit indices and edges represent the input connectivity of the associated unitary. Each motif type corresponds to a different method of edge generation. The cycle iterates through the list of nodes when creating edges, while the pivot connects one or more fixed nodes to all remaining nodes. The mask motif makes nodes unavailable to subsequent motifs, while its counterpart the unmask makes unavailable nodes available again. The mask motif can be associated with non-unitary operations, specifically operations conditioned on measurement outcomes, however, we have not used this functionality in the experiments presented here.
The specific behaviour of each motif is controlled by several properties that influence edge generation. For example, the global pattern property of the pivot is used to control which nodes are fixed in the one-to-many edge generation pattern. Once the number of nodes is fixed, the graph’s edges are generated and the associated unitary is repeatedly connected to each edge, resulting in a quantum circuit.
@@
A hierarchical representation of a large architecture is an effective way to reduce search complexity and leads to better search efficiency Liu et al. (2018); Ru et al. (2020); Christoforidis et al. (2023). By limiting the search space to a small number of motifs, that in turn depend on a limited number of properties, we can significantly reduce the size of the search space compared to a gate-sequence circuit encoding. Importantly, we are able to reduce the search space complexity without sacrificing expressivity. Algorithms are inherently scalable and our DSL captures this property. Further, our DSL biases the search space toward structured circuits which are likely to include the structured operations necessary to build algorithms. In this way, we were able to automatically design large circuit structures in reasonable times, as demonstrated in Fig. 4 and Table 1.
Our DSL simplifies the search strategy compared to approaches that use a gate-based circuit representation. Using the DSL, it is no longer necessary to construct a set of rules to govern gate placements within the circuit during search Rattew et al. (2020). Instead, we define a configuration space, that is a set of available discrete values for each property of the motifs. This in turn determines the set of possible algorithms that can be constructed.

### v3

# Automated Quantum Algorithm Design using a Domain-Specific Language

@@ I Introduction
Quantum computers promise exponential speed-ups, yet there are only a handful of classes of quantum algorithms that realise this. Shor shor2003haven provides us with two possible explanations. Firstly, quantum computers operate so differently from classical computers that our existing techniques for algorithm design and our intuitions about computation may no longer be effective. Secondly, it is possible that only a limited number of such quantum algorithms exist and that many or all of them may have already been discovered.
@@
Given the steadily growing Quantum Algorithm Zoo  [^1] https://quantumalgorithmzoo.org/ we might retain some optimism with regards to the latter point. Nevertheless, developing novel protocols—especially those that generate new classes of quantum algorithms, of which Shor’s factorising algorithm shor1999polynomial is an example—is clearly a difficult task. Computational methods could offer a way to bridge the intuition gap and facilitate the discovery process. We propose such a method that uses a bottom-up approach, where a collection of well-constructed examples is translated into an algorithm.
@@
Our approach follows the methodology of neural architecture search (NAS) where the task of automated design is divided into three parts, search space, search strategy and performance estimation strategy elsken2019neural; wistuba2019survey; liu2021survey; white2023neural. Neural architecture search strategies are widely used in machine learning to improve performance in computer vision, natural language processing and time series forecasting liu2021survey. More recently, NAS methods have also been applied in the context of variational quantum algorithms zhang2022differentiable, for a recent survey see martyniuk2024quantum.
@@
Quantum algorithms, which are associated with highly structured circuits, tend to be modular and to contain repeating patterns. To exploit these facts we use a domain-specific language (DSL)—specific syntax and rules for compactly expressing quantum circuits via modular building blocks, implemented as an open source Python package lourens2023hierarchical; lourensHierarQcalQuantumCircuit2024; lourensHierarqcalGithub2024. In this DSL, quantum circuits are abstracted beyond the usual gate-sequence description and scale automatically with problem size. This allows us to optimise the structure of the algorithm, as opposed to the quantum circuit at the level of individual gates. Previous work on automated quantum algorithm design almuqbil2024discovery; lin2020quantum; wan2018learning; potovcek2018multi; bang2014strategy; bang2014genetic; hutsell2007applying; surkan2002evolutionary was limited to gate-based circuit implementations of algorithms for small circuit sizes, usually 22-88 qubits. In contrast, our learnt algorithms automatically generalise to arbitrary problem size which has not been demonstrated previously. Remarkably, this was achieved by using training examples related to just three small circuit sizes, namely $2$, 33 and $4$ qubits or $3$, 44 and $5$ qubits.
@@
Our method is compatible with any gradient-free optimisation procedure, specifically, we use evolutionary search de1989genetic; schwefel1995contemporary; rechenberg1973evolutionsstrategie. The performance of candidate algorithms is measured by how closely their behaviour matches the examples provided in the training set. In addition, the evaluation takes into account the uniformity of the performance across all examples and the complexity of the candidate algorithm. We demonstrate our methodology by rediscovering three well-known quantum algorithms, the quantum Fourier transform, the Deutsch-Jozsa algorithm and Grover’s search. This shows that our method works for oracle-based problems and variational circuits. Our approach also has the necessary functionality for designing both measurement-based algorithms and ansatzes for variational quantum circuits lourens2023hierarchical; lourens2025generatinggeneralisedgroundstateansatzes.
@@
grover = ( Oracle + QCycle(H) + QCycle(X) + QPivot(H, ”*1”) + QPivot(MCX, ”*1”) + QPivot(H, ”*1”) + QCycle(X) + QCycle(H) ) * r
[…]
Figure 1: Two equivalent $n$-qubit representations of three learnt algorithms, namely the rediscovered implementations of the Deutsch-Jozsa algorithm (top), Grover’s search (middle) and the quantum Fourier transform (bottom). On the left are DSL descriptions output by our search method, while our experiments found many variations, one example per task is shown. On the right are the corresponding circuit representations that match those typically presented in the literature. The cycle (QCycle—) and pivot (QPivot—) motifs encode two different gate placement patterns, as can be seen by comparing the left- and right-hand sides. The mask (QMask—) motif makes qubits unavailable. The strings ”1*”— and ”*1”— are examples of patterns used to control the behaviour of the mask and pivot motifs, and represent the selection of the first and last qubits, respectively. The parameter r— controls the number of times the structure is repeated, where the relationship to the circuit size can be inferred from training.
@@ II Results
Using our domain-specific language combined with evolutionary search we were able to rediscover three well-known quantum algorithms, the quantum Fourier transform (QFT), the Deutsch-Jozsa algorithm and Grover’s search. Figure 1 shows two equivalent $n$-qubit representations of three automatically designed algorithms, corresponding to each of the learning tasks. On the left-hand side of Fig. 1 are the leant algorithms expressed in the DSL. On the right-hand side are the corresponding quantum circuit representations, where each circuit matches the common, hardware-agnostic, implementation of each algorithm nielsen2001quantum. Fig. 1 illustrates how the DSL compactly describes the general $n$-qubit implementation of the algorithms. This allows us to immediately interpret the algorithm structure, as well as identify symmetries and patterns. For example, the symmetric structure of the diffusion term in Grover’s search is immediately identifiable.
An evolutionary algorithm is used to explore the search space where new candidate algorithms are generated by changing or combining the fittest candidates from a population of algorithms. Remarkably, we found that to converge to the general algorithm, it was sufficient to evaluate each candidate on three small circuit sizes, namely $2$, 33 and $4$ qubits for the Deutsch-Jozsa algorithm and QFT, and $3$, 44 and $5$ qubits for Grover’s search. A candidate algorithm has a high fitness if it executes the target algorithm correctly and has low complexity, discussed further in section IV.3. Furthermore, we did not tune the hyper-parameters of the evolutionary search. These were kept as consistent as possible across all experiments. That the same hyper-parameters were good enough for all three tasks indicates that our search strategy is stable and reliable.
@@
The right panel of Fig. 2 shows the changing composition of the population. The slope of the contour lines indicates the rate of displacement of the poorest performing candidates from the top $200$ algorithms. The steeper the slope of a contour line, the slower poorer-performing candidates are displaced. The general trend is that the gradient of each contour line decreases over time as the current best algorithm, its duplicates, and related variants propagate rapidly and take up an increasing proportion of the population.
Fig. 3 shows the evolutionary path, represented as a directed graph, that led to Grover’s search. The graph contains a roughly equal distribution of mutation types, represented by different edges. This indicates that all four mutation types played an important role in producing the final algorithm, which was the case for all experiments across all three tasks. In the literature, it has been suggested that mutations that combine two candidates, in our case the ‘chop-and-join’ mutation, do not lead to better-performing architectures white2023neural; rattew_domain-agnostic_2020. Our results indicate that in the context of algorithm design, this is not the case. A possible explanation is that we break each architecture into two pieces before recombining them, rather than simply appending algorithms head-to-tail. This provides an opportunity for useful sections of each algorithm to be combined. Dividing each algorithm into two randomly determined pieces before recombination also ensures that the size of candidates does not double with each application of this mutation. Furthermore, random division ensures diversity in the size of candidates.
@@
We demonstrate the robustness of our method by investigating the effect of increasing search space size. The DSL that was used to represent our search space is implemented as an open-source Python package called Hierarqcal lourens2023hierarchical; lourens2025generatinggeneralisedgroundstateansatzes. See Section IV.1 and Fig. 1 for an illustration of the basic functionality of Hierarqcal. In this framework, algorithms are built up of motifs that encode different gate placement patterns. Each motif has several properties, each of which can take on a possibly infinite number of discrete values. Motif properties include mapping, step, stride, offset, boundary condition, global and merge within patterns, weight sharing and edge order. Selecting appropriate sets of discrete values for each of these properties determines the configuration space, which in turn determines the set of possible algorithms that can be constructed, i.e. the search space.
@@
In Fig. 4 we see that the number of candidates evaluated in each experiment was a small fraction of the size of the search space, indicating the efficiency of our evolutionary strategy. The number of evaluations required grows with the size of the search space, though favourably at a decreasing rate. A total of $50$ experiments were performed for each combination of task and search space size, and successful runs are shown in Fig. 4. An experiment is considered successful if it is completed within a certain runtime limit and if certain stopping conditions are met, which we describe in detail in Section IV.3. Table 1 shows the approximate success rates for each task and notes the runtime limits in core processing unit (CPU) hours. Taking the success rates into account, our method is robust and can plausibly be run on a personal computer – $8$ CPUs – with a $90\%$ chance of success within two days in most cases.
@@
When parametrised unitary gates are used, as was the case for the QFT algorithm, runtimes are longer since the parameters of these gates have to be trained during evaluation. We note that, by inspecting the learnt gate parameters it is possible to infer the general relationship between parameters, which may also depend on the circuit sizes. Similarly for all tasks, the relationship between the circuit size and number of repetitions, r in Fig. 1, can also be inferred in this way. For example, for QFT we saw that the learnt number of repetitions was equal to the circuit size.
@@ III Discussion
Our results demonstrate the utility of our approach as a tool for quantum algorithm design. We successfully rediscover the quantum Fourier transform (QFT), the Deutsch-Jozsa algorithm, and Grover’s search, validating our approach and outlining how it could be used to discover a new algorithm given a well-posed task. Beyond quantum algorithms, our approach is also applicable to the design of variational Ansätze for quantum many-body states, as demonstrated in lourens2025generatinggeneralisedgroundstateansatzes. In lourens2025generatinggeneralisedgroundstateansatzes, the same DSL allows for the discovery of new interpretable wavefunctions that remain analytically tractable. Applied to ground-state problems, the method autonomously rediscovers a mean-field description and extends it to incorporate correlations, producing wavefunctions that extrapolate to arbitrary system size without modification.
Central to our method is a domain-specific language, which captures the inherent scalability of algorithms. Further, the DSL biases the algorithms in the search space to be structured, capturing modularity and repeating patterns. As a result, the learnt algorithms are immediately interpretable and analysable. Notably, convergence to the general algorithm requires evaluation on only three small circuit sizes, containing at most $5$ qubits, which significantly reduces the computational cost.
@@
Furthermore, the DSL contains a small number of building blocks that in turn depend on a limited number of properties. Therefore, we can greatly reduce the size of the search space compared to using a gate-based circuit description. Importantly, we are able to reduce the search space size without sacrificing expressivity, since the DSL captures the essential characteristics of algorithms, namely structure and scalability.
@@
By using a bottom-up approach, we reformulate the task of algorithm design as a task of creating a set of well-constructed training examples. Our performance estimation strategy then evaluates candidate algorithms based on their ability to generalise across these training examples. We prioritise simpler algorithms by incorporating a complexity penalty while simultaneously rewarding algorithms that perform uniformly across training examples. This, combined with the attributes mentioned above, enables us to automatically design the $n$-qubit realisation of a quantum algorithm.
@@ IV Methods
Much of our success is due to the use of a domain-specific language. It allows us to learn the general algorithm structure rather than a specific unitary implementation of the algorithm for some number of qubits. The DSL was adapted from its initial formalism detailed in lourens2023hierarchical, where many of the core ideas were introduced. For a recent overview, see the blog post lourensHierarQcalQuantumCircuit2024, and for the associated Python package Hierarqcal, see the GitHub repository lourensHierarqcalGithub2024.
The original inspiration for the DSL came from the field of neural architecture search liu_hierarchical_2018. In line with NAS we divide our automated algorithm design approach into three components: search space, search strategy, and performance estimation elsken2019neural; wistuba2019survey; liu2021survey. The DSL is used to specify the search space, see Section IV.1. Owing to the modular nature of our DSL, evolutionary search is a natural choice for search strategy, detailed in Section IV.2. The performance estimation method used to evaluate candidate algorithms is described in Section IV.3. Finally, in Section IV.4, we provide an overview of each design task, with a particular focus on problem setup.
@@ IV.1 Search space
The motifs are the building blocks of our DSL and are used to generate hypergraphs. We denote a hypergraph by $\mathcal{G}=(Q,E)$, where the sequences $Q$ and $E$ list qubit indices and edges, respectively. Each edge $e\in E$ encodes the connectivity of an associated unitary operation $U_{e}$. For example, a $3$-qubit unitary gate $U_{3,1,5}$ acts on qubits three, one and five, and is represented by the edge $e=(3,1,5)$ in the hypergraph. A candidate algorithm $\mathcal{A}=(M_{1},\dots,M_{p})$ is a sequence of $p$ motifs. For example, the QFT, Fig. 1 left, is a sequence of three motifs, two consecutive pivots followed by a mask, repeated $r$ times. A candidate algorithm $\mathcal{A}$ is instantiated by specifying a list of qubit indices $Q$. Each motif $M_{m}$ generates the hypergraph $\mathcal{G}_{m}=(Q,E_{m})$ with associated unitary operation $U^{(m)}$. The list of qubit indices $Q$ does not carry a subscript $m$ since the edges of all hypergraphs encode unitary operations acting on the same quantum circuit. The quantum circuit, with qubit indices $Q$, corresponding to a candidate algorithm $\mathcal{A}$ is given by
$U\left(Q,\mathcal{A}=\left(M_{1},\dots,M_{p}\right)\right)$ $=\mathop{\prod}\limits_{m\in[1,p]}^{\leftarrow}\ \mathop{\prod}\limits_{e\in E_{m}}^{\leftarrow}U^{(m)}_{e}$, (1)
where the product of unitaries is ordered with indices increasing from right to left, as indicated by the arrows. In Eq. (1), we see that the sequence of edges $E_{m}$ determines the gate placement patterns within the circuit. In turn, each motif $M_{m}$ generates the sequence of edges $E_{m}$ based on a rule that captures size scaling. In this way, the algorithm $\mathcal{A}$ is built without having to specify a system size. Eq. (1) enables the analytical analysis of the resulting circuit, and allows us to verify whether the generalisation of the algorithm to any problem size is correctly captured.
Each motif type corresponds to a different method of edge generation. The cycle iterates through the list of nodes when creating edges, while the pivot connects one or more fixed nodes to all remaining nodes. The mask motif makes nodes unavailable to subsequent motifs, while its counterpart the unmask makes unavailable nodes available again. The specific behaviour of each motif type is controlled by several properties that influence edge generation. For example, the global pattern property of the pivot is used to control which nodes are fixed in the one-to-many edge generation pattern. Appendix A provides further explanation of the DSL by example.
@@
A hierarchical representation of a large architecture is an effective way to reduce search complexity and leads to better search efficiency liu_hierarchical_2018; ru_neural_2020; christoforidis_novel_2023. By limiting the search space to a small number of motifs, that in turn depend on a limited number of properties, we can significantly reduce the size of the search space compared to a gate-sequence circuit encoding. Importantly, we are able to reduce the search space complexity without sacrificing expressivity. Algorithms are inherently scalable and our DSL captures this property. Further, our DSL biases the search space toward structured circuits which are likely to include the structured operations necessary to build algorithms. In this way, we were able to automatically design large circuit structures in reasonable times, as demonstrated in Fig. 4 and Table 1.
Our DSL simplifies the search strategy compared to approaches that use a gate-based circuit representation. Using the DSL, it is no longer necessary to construct a set of rules to govern gate placements within the circuit during search rattew_domain-agnostic_2020. Instead, we define a configuration space, that is a set of available discrete values for each property of the motifs. This in turn determines the set of possible algorithms that can be constructed. For example, for Grover’s search algorithm and the largest configuration space, labelled Huge in Fig. 4, the number of possible unique motifs is $646$, where the oracle is also counted as a motif. This results in a search space size of approximately $3\times 10^{22}$, where as discussed in Sec. II the search space size is the number of algorithms that would be evaluated in a worst-case-scenario brute-force search  [^2] A count of all the possible candidate algorithms consisting of up to $p$ motifs is given by $646+646^{2}+\cdots+646^{p}={646(646^{p}-1)}/({646-1})$. The minimum number of motifs required to represent Grover’s search algorithm is eight. . In contrast, for an equivalent gate-based search method, the search space size scales exponentially with circuit size. In our method, the search space size is independent of the number of qubits.
In almuqbil2024discovery the authors also simplified the search space by using a vector to encode the sequence of gates to be applied, where it was assumed that each gate in the sequence is applied to all qubits. In the case of two-qubit gates, these are applied from each qubit to the next without looping back, which is equivalent to a cycle with open boundary conditions in our DSL. Using this representation, the authors were able to rediscover Grover’s search algorithm for circuits containing 33-88 qubits. In their vector-based representation, the assumption that a gate acts on each qubit in the circuit was motivated by the permutation symmetry of Grover’s search and therefore does not generalise to other algorithms that do not have the same permutation symmetry, including QFT and the Deutsch-Jozsa algorithm.